OpenStudy (anonymous):

Find the diameter of a circle whose area is 254.34 square inches.

OpenStudy (klimenkov):

\(\text{Area of a circle}=\pi\cdot (\frac{\text{Diameter}}2)^2\)

OpenStudy (anonymous):

radius=8.99 inch rounded 9 inch